# Projection-free computation of robust controllable sets with constrained zonotopes

### v1

@@ 8 Conclusion
We presented novel theory and algorithms to approximate the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend. For broad classes of subtrahends, our approach admits closed-form expressions for the inner-approximation. Then we used these algorithms for a tractable and scalable computation of an inner-(and outer-)approximation of the robust controllable set for discrete-time linear systems with additive disturbance sets that are symmetric, convex, and compact and subject to linear state and input constraints. We showed by numerical simulations that the proposed approach provides non-trivial inner-approximations of the RC sets with significantly shorter computation times than the previously published methods.
A current limitation of the proposed approach is that the subtrahend must be a symmetric set, which prevents us from stating a generic closed-form expression for the Pontryagin difference with a constrained zonotopic minuend, similar to (8) for a polytopic minuend. Our future work will investigate how to relax this assumption, and develop methods to apply the proposed algorithms in systems with low computational resources.
@@ Appendix A Proofs
A.1 Proof of Proposition 1
Assume for contradiction that, for some representation of $\mathcal{C}$, some of the rows in $G_{C}$ are linearly dependent. Then, there exists a vector $\alpha\in\mathbb{R}^{n},\ \alpha\neq 0$ such that $\alpha^{\top}G_{C}=0$. From (3), for every $x\in\mathcal{C}$, there exists $\xi\in\mathcal{B}_{\infty}(A_{C},b_{C})$ such that $G_{C}\xi=x-c_{C}$. Consequently, for any $x\in\mathcal{C}$, $\alpha^{\top}(x-c_{C})=\alpha^{\top}G_{C}\xi=0$, implying that $\mathcal{C}\subset\{x:\alpha^{\top}(x-c_{C})=0\}$. Thus, the affine dimension of $\mathcal{C}$ is smaller than $n$, a contradiction since $\mathcal{C}$ is full-dimensional. Thus, $G_{C}$ has full row rank.
Let $[A_{C}^{\prime},b_{C}^{\prime}]$ be a set of linearly independent $M_{C}^{\prime}\leq M_{C}$ rows of the matrix $[A_{C},b_{C}]$ such that $\operatorname{rank}([A_{C}^{\prime},b_{C}^{\prime}])=\operatorname{rank}([A_{C},b_{C}])$. We show that $\mathcal{C}=(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$. Without loss of generality, assume that $[A_{C}^{\prime},b_{C}^{\prime}]$ are the first $M_{C}^{\prime}$ rows of $[A_{C},b_{C}]$. By the choice of $[A_{C}^{\prime},b_{C}^{\prime}]$, there exists $E\in\mathbb{R}^{(M_{C}-M_{C}^{\prime})\times M_{C}^{\prime}}$ such that $[A_{C},b_{C}]=[I_{M_{C}^{\prime}};E][A_{C}^{\prime},b_{C}^{\prime}]$. Since the matrix $[I_{M_{C}^{\prime}};E]$ has full column rank, $[I_{M_{C}^{\prime}};E]y=0$ if and only if $y=0$. Consequently, the solution set $\{\xi:A_{C}\xi=b_{C}\}=\{\xi:A_{C}^{\prime}\xi=b_{C}^{\prime}\}$, since for any $\xi$ such that $A_{C}\xi=b_{C}$, $[A_{C},b_{C}][\xi;-1]=0\Leftrightarrow[I_{M_{C}^{\prime}};E][A_{C}^{\prime},b_{C}^{\prime}][\xi;-1]=0\Leftrightarrow[A_{C}^{\prime},b_{C}^{\prime}][\xi;-1]=0$. Thus, $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})=(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$.
Next, we observe that $A_{C}^{\prime}$ also has a full row rank of $M_{C}^{\prime}$, using the following observations: 1) the matrix $[A_{C}^{\prime},b_{C}^{\prime}]$ has full row rank of $M_{C}^{\prime}$, 2) the set $\{\xi:A_{C}^{\prime}\xi=b_{C}^{\prime}\}$ is non-empty, and 3) $b_{C}^{\prime}$ lies in the column space of $A_{C}^{\prime}$.
Finally, we show that the matrix $[G_{C};A_{C}^{\prime}]$ has full row rank to conclude that the constrained zonotope $(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$ is a MinRow representation of $\mathcal{C}$. Assume, for contradiction, that the matrix $[G_{C};A_{C}^{\prime}]$ has linearly dependent rows. In other words, there exists $\beta_{1}\in\mathbb{R}^{n},\ \beta_{1}\neq 0$ and $\beta_{2}\in\mathbb{R}^{M_{C}^{\prime}},\ \beta_{2}\neq 0$ s.t.,
@@ A.1 Proof of Proposition 1
We know $\beta_{1}^{\top}G_{C}\neq 0$ and $\beta_{2}^{\top}A_{C}^{\prime}\neq 0$, since all rows of $G_{C}$ are linearly independent, and all rows of $A_{C}^{\prime}$ are linearly independent. From (1b), for all $x\in\mathcal{C}=(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$, there exists $\xi\in\mathbb{R}^{N_{C}}$ s.t.,
$[G_{C};A_{C}^{\prime}]\xi=[x-c_{C};b_{C}^{\prime}]\text{ and }\|\xi\|_{\infty}\leq 1,.$ (80)
By (79) and (80), $\mathcal{C}\subset\{x:\beta_{1}^{\top}(x-c_{C})+\beta_{2}^{\top}b_{C}^{\prime}=0\}$, which is a contradiction since $\mathcal{C}$ is full-dimensional. Thus, all rows of $[G_{C};A_{C}^{\prime}]$ are linearly independent. ∎
A.2 Proof of Proposition 2
@@
$\mathcal{C}$ =$=\left\{x:\inf_{\begin{subarray}{c}\xi\in\mathbb{R}^{N_{C}}\\ [G_{C};A_{C}]\xi=[x-c_{C};b_{C}]\end{subarray}}{\|\xi\|}_{\infty}\leq 1\right\}$ (83)
=$=\left\{x:\sup_{\begin{subarray}{c}\nu\in\mathbb{R}^{n+M_{C}}\\ {\|[G_{C};A_{C}]^{\top}\nu\|}_{1}\leq 1\end{subarray}}\nu^{\top}[x-c_{C};b_{C}]\leq 1\right\}.$ (86)
For every $i\in\mathbb{N}_{[1:N_{C}]}$, define $\nu_{i}^{\ast}$ as the minimum-norm solution to $[G_{C};A_{C}]^{\top}\nu=e_{i}$. Under Assumption 1, $[G_{C};A_{C}]^{\top}$ has full column rank by Proposition 1, and $\nu_{i}^{\ast}$ is available in closed-form, where
@@
for every $i\in\mathbb{N}_{[1:N_{C}]}$. Step 3 of Algorithm 2 (see (42)) rescales $\nu_{i}^{\ast}$ to ensure that $[G_{C};A_{C}]^{\top}\nu_{i}^{\ast}$ has a unit $\ell_{1}$-norm to obtain $v_{i}^{\top}$. Thus, ±$\pm v_{i}^{\top}$ are feasible for (86) for every $i\in\mathbb{N}_{[1:N_{C}]}$, and using (86),
$\mathcal{C}$$\subseteq\underbrace{\left\{{x:{\left\|V^{\top}[x-c_{C};b_{C}]\right\|}_{\infty}\leq 1}\right\}}_{\mathcal{P}\triangleq\left\{x:Hx\leq k\right\}},$ (87)
with $V=[\nu^{\ast}_{1};\nu^{\ast}_{2};\ldots;\nu^{\ast}_{N_{C}}]$. Since $\mathcal{P}$ is a finite intersection of $2N_{C}$ hyperplanes in $\mathbb{R}^{n}$, $\mathcal{P}$ is a convex polyhedron with $H\in\mathbb{R}^{(2N_{C})\times n}$ and $k\in\mathbb{R}^{2N_{C}}$, as defined in Steps 4 and 5 of Algorithm 2. ∎
A.3 Proof of Proposition 4
Since $[G_{C};A_{C}]$ is invertible under Assumption 2, the optimal solution to (83) in Appendix A.2 has a closed-form expression $\xi^{\ast}={[G_{C};A_{C}]}^{-1}[x-c_{C};b_{C}]$. Consequently,
$\mathcal{C}$ =$=\{x:{\|{[G_{C};A_{C}]}^{-1}[x-c_{C};b_{C}]\|}_{\infty}\leq 1\}$ (88)
=$=\left\{x\middle|\forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}{[G_{C};A_{C}]}^{-1}[x-c_{C};b_{C}]\leq 1\right\}.$ (89)
Here, $v_{i}$ defined in Step 3 of Algorithm 2 also simplifies to $e_{i}^{\top}{[G_{C};A_{C}]}^{-1}$ when $[G_{C};A_{C}]$ is invertible [20, Sec. 11.5]. Thus, Algorithm 2 returns a convex polytope $\mathcal{P}$ that coincides with (89) (thereby, equal to $\mathcal{C}$) under Assumption 2. ∎
A.4 Proof of Proposition 5
Exactness of $\mathcal{M}^{+}$) Under Assumption 2, Algorithm 2 computes an exact polytopic representation of $\mathcal{C}$. Since the rest of steps of Algorithm 3 are exact, $\mathcal{M}^{+}=\mathcal{C}\ominus\mathcal{S}$.
Exactness of $\mathcal{M}^{-}$) Since $[G_{C};A_{C}]$ is invertible under Assumption 2, $\Gamma$ prescribed by Theorem 2 is
@@ A.4 Proof of Proposition 5
by [20, Sec. 11.5]. We show that $\mathcal{M}$ is an affine transformation of $\mathcal{C}_{L}$ defined in the proof of Theorem 1. Specifically,
$\mathcal{M}$ $=\mathcal{C}\ominus\mathcal{S}=\{x\ |\ \forall s\in\mathcal{S}_{0},\ x+c_{S}+s\in\mathcal{C}\}$
@@
=$=\left\{x\middle|\ \forall s\in\mathcal{S}_{0},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}{[G_{C};A_{C}]}^{-1}[x+s+c_{S}-c_{C};b_{C}]\leq 1\right\}$ (91)
@@
=$=\left\{x\ \middle|\ \forall s\in\mathcal{S}_{0},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\left({\Gamma x+\Gamma s+{[G_{C};A_{C}]}^{-1}[-c_{M^{-}};b_{C}]}\right)\leq 1\right\}$ (92)
@@
=$=\left\{x\ \middle|\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\left({\Gamma x+{[G_{C};A_{C}]}^{-1}[-c_{M^{-}};b_{C}]}\right)\leq 1-\rho_{\mathcal{S}_{0}}(\Gamma^{\top}e_{i})\ \right\}$ (93)
@@
=$=\left\{x\ \middle|\ \xi\triangleq\Gamma x+{[G_{C};A_{C}]}^{-1}[-c_{M^{-}};b_{C}],\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\xi\leq D_{ii}\ \right\}$ (94)
@@
=$=\left\{x\ \middle|\ G_{C}\xi+c_{M^{-}}=x,\ A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\xi\leq D_{ii}\ \right\}$ (95)
=$=\left\{G_{C}\xi+c_{M^{-}}\ \middle|\ A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ |e_{i}^{\top}\xi|\leq D_{ii}\right\}=G_{C}\mathcal{C}_{L}+c_{M^{-}}=\mathcal{M}^{-}.$ (96)
@@
Here, (91) follows from (6d) and (89) in Appendix A.3, (92) follows from (90), (93) follows from encoding the condition for all $s\in\mathcal{S}_{0}$ using the support function of $\mathcal{S}_{0}$, (94) follows from the definition of $D_{ii}$ in (26), (95) follows from expressing $\xi$ as a solution of linear equations $[G_{C};A_{C}]\xi=[x-c_{M^{-}};b_{C}]$ with invertible $[G_{C};A_{C}]$, and (96) follows from the definition of $\mathcal{C}_{L}$ in (35) (see the proof of Theorem 1). Consequently, under Assumption 2, Theorem 2 provides an exact characterization of $\mathcal{C}\ominus\mathcal{S}$. ∎
A.5 Proof of Proposition 6
From [23, Thm. 2.1(iii)],
$(\mathcal{U}\ominus\mathcal{V})\oplus\mathcal{V}\subseteq\mathcal{U}.$ (97)
@@ A.5 Proof of Proposition 6
Let a failure occur at some time $s\in\mathbb{N}$. Using (61) and (97), for some $i\in\mathbb{N}_{[1:N_{D}]}$, $x_{s}=\hat{z}_{s|s}\in\mathcal{K}(s,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})\ominus\mathcal{E}_{\text{off-nom}}$, which implies that $\hat{z}_{s|s}\in\mathcal{K}(s,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})$. From Definition 2, for every time step $k\in\mathbb{N}_{[s:s+T_{\text{safe}}]}$, there exists a feedback law $\pi_{k}\in\Pi$ such that $u_{k|s}=\pi_{k}(\hat{z}_{k|s})\in\mathcal{U}_{\text{off-nom}}$ steers the state estimate according to the dynamics (60) to satisfy $\hat{z}_{k|s}\in\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$, irrespective of the disturbance $\phi_{k|s}\in\Phi_{k}$. By (61) and (97), $\hat{z}_{k|s}\in\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$ implies that the true post-failure state $z_{k|s}\in\mathcal{H}_{i}^{\complement}$ (and thereby, $z_{k|s}\not\in\mathcal{D}$) for all $k\in\mathbb{N}_{[s:s+T_{\text{safe}}]}$, which completes the proof.

### v2

@@ 8 Conclusion
We presented novel theory and algorithms to approximate the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend. For broad classes of subtrahends, our approach admits closed-form expressions for the inner-approximation. We use these algorithms for a tractable and scalable computation of an inner-(and outer-)approximation of the robust controllable set for discrete-time linear systems with additive disturbance sets that are symmetric, convex, and compact and subject to linear state and input constraints. We showed by numerical simulations that the proposed approach provides non-trivial inner-approximations of the RC sets with significantly shorter computation times than the previously published methods.
@@ Appendix A Proofs
A.1 Proof of Prop. 1
Proof of 1) Assume for contradiction that, $G_{C}$ does not have full row rank for some full-dimensional constrained zonotope $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$. Then, there exists a vector $\alpha_{G_{C}}\in\mathbb{R}^{n},\ \alpha_{G_{C}}\neq 0$ such that $\alpha_{G_{C}}^{\top}G_{C}=0$. From (2), for every $x\in\mathcal{C}$, there exists $\xi\in\mathcal{B}_{\infty}(A_{C},b_{C})$ such that $G_{C}\xi=x-c_{C}$. Consequently, $\mathcal{C}\subset\{x\mid\alpha_{G_{C}}^{\top}x=\alpha_{G_{C}}^{\top}c_{C}\}$ since $\alpha_{G_{C}}^{\top}(x-c_{C})=\alpha_{G_{C}}^{\top}G_{C}\xi=0$ for every $x\in\mathcal{C}$. In other words, the affine dimension of $\mathcal{C}$ is smaller than $n$. However, this is a contradiction since $\mathcal{C}$ is full-dimensional. Thus, $G_{C}$ has full row rank for every full-dimensional $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$.
Proof of 3) We now show that Algo. 1 converts any full-dimensional $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$ into a MinRow representation. From 1), $G_{C}^{\prime}=G_{C}$ has full row rank. Consider $A_{C}^{\prime},b_{C}^{\prime}$ in Step 4 of Algo. 1 with $[A_{C}^{\prime},b_{C}^{\prime}]\in\mathbb{R}^{M_{C}^{\prime}\times(N_{C}+1)}$. Then, $M_{C}^{\prime}=\operatorname{rank}([A_{C}^{\prime},b_{C}^{\prime}])=\operatorname{rank}([A_{C},b_{C}])\leq M_{C}$. It suffices to show that $\mathcal{C}=(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$ and $[G_{C};A_{C}^{\prime}]$ has full row rank.
Without loss of generality, assume that $A_{C}^{\prime},b_{C}^{\prime}$ are the first $M_{C}^{\prime}$ rows of $[A_{C},b_{C}]$. Since $\operatorname{rank}([A_{C},b_{C}])=\operatorname{rank}([A_{C}^{\prime},b_{C}^{\prime}])$, every row of $[A_{C},b_{C}]$ is a linear combination of the rows of $[A_{C}^{\prime},b_{C}^{\prime}]$. In other words, there exists $E\in\mathbb{R}^{({M_{C}}-{M_{C}}^{\prime})\times{M_{C}}^{\prime}}$ such that $[A_{C},b_{C}]=[I_{{M_{C}}^{\prime}};E][A_{C}^{\prime},b_{C}^{\prime}]$. The matrix $[I_{{M_{C}}^{\prime}};E]$ has full column rank implying that $[I_{{M_{C}}^{\prime}};E]y=0$ if and only if $y=0$ [21, Ch. 11]. Thus, $\{\xi\mid A_{C}\xi=b_{C}\}=\{\xi\mid A_{C}^{\prime}\xi=b_{C}^{\prime}\}$, since for any $\xi$ such that $A_{C}\xi=b_{C}$, $[A_{C},b_{C}][\xi;-1]=0\Leftrightarrow[I_{{M_{C}}^{\prime}};E][A_{C}^{\prime},b_{C}^{\prime}][\xi;-1]=0\Leftrightarrow[A_{C}^{\prime},b_{C}^{\prime}][\xi;-1]=0$. Therefore, $\mathcal{B}_{\infty}(A_{C},b_{C})=\mathcal{B}_{\infty}(A_{C}^{\prime},b_{C}^{\prime})$ and $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})=(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$.
Since full-dimensional constrained zonotopes are non-empty, $\{\xi\mid A_{C}^{\prime}\xi=b_{C}^{\prime}\}$ is non-empty and $\operatorname{rank}(A_{C}^{\prime})=\operatorname{rank}([A_{C}^{\prime},b_{C}^{\prime}])=M_{C}^{\prime}$. Assume, for contradiction, that the matrix $[G_{C};A_{C}^{\prime}]$ has linearly dependent rows. In other words, there exists $\beta_{1}\in\mathbb{R}^{n},\ \beta_{1}\neq 0$ and $\beta_{2}\in\mathbb{R}^{{M_{C}}^{\prime}},\ \beta_{2}\neq 0$ s.t.,
@@
We know $\beta_{1}^{\top}G_{C}\neq 0$ and $\beta_{2}^{\top}A_{C}^{\prime}\neq 0$, since rows of $G_{C}$ and $A_{C}^{\prime}$ are linearly independent among themselves. From (1b), for all $x\in\mathcal{C}$, there exists $\xi\in\mathbb{R}^{{N_{C}}}$ s.t.,
$[G_{C};A_{C}^{\prime}]\xi=[x-c_{C};b_{C}^{\prime}]\text{ and }\|\xi\|_{\infty}\leq 1.$ (87)
By (86) and (87), $\mathcal{C}\subset\{x\mid\beta_{1}^{\top}(x-c_{C})+\beta_{2}^{\top}b_{C}^{\prime}=0\}$, which is a contradiction since $\mathcal{C}$ is full-dimensional. Thus, all rows of $[G_{C};A_{C}^{\prime}]$ are linearly independent, and $(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$ is a MinRow representation of $\mathcal{C}$.
Proof of 2) Since constrained zonotopes are a representation of polytopes [5, Thm. 1], the proof of 3) also shows that every full-dimensional polytope admits a MinRow representation. We now show that the reverse is also true, i.e., any non-empty polytope $\mathcal{C}$ in MinRow representation $(G_{C},c_{C},A_{C},b_{C})$ is full-dimensional. Assume, for contradiction, that $\mathcal{C}$ is not full-dimensional. Then, there exists an affine set characterized by $\alpha\in\mathbb{R}^{n},\alpha\neq 0$ and $\beta\in\mathbb{R}$ such that $\mathcal{C}\subset\{x\mid\alpha^{\top}x=\beta\}$. Since $\mathcal{C}$ is non-empty, $\mathcal{C}=\mathcal{C}\cap\{x\mid\alpha^{\top}x=\beta\}$. From [6], $\mathcal{C}=(G_{C},c_{C},[A_{C};\alpha^{\top}G_{C}],[b_{C};\beta-\alpha^{\top}c_{C}])=(G_{C},c_{C},A_{C},b_{C})$. Consequently, there exists some $\gamma\in\mathbb{R}^{{M_{C}}},\gamma\neq 0$ such that $\alpha^{\top}G_{C}=\gamma^{\top}A_{C}$ and $\beta-\alpha^{\top}c_{C}=\gamma^{\top}b_{C}$, a contradiction since $[G_{C};A_{C}]$ has full row rank. Thus, $\mathcal{C}$ is full-dimensional, which completes the proof.
Proof of 4) Algo. 2 follows the steps of [5, Thm. 1] with the zonotope $\mathcal{Z}$ in Step 1 defined as a rectangular outer-approximation with $\mathcal{C}\subseteq\{x\in\mathbb{R}^{n}\mid l\leq x\leq u\}$, and $\sigma\in\mathbb{R}^{{L_{C}}}$ defined in Step 2 using (4) such that $\mathcal{P}=\{x\in\mathcal{Z}\mid\sigma\leq Hx\leq k\}\subset\mathbb{R}^{n}$. Thus, $(G_{C},c_{C},A_{C},b_{C})$ defined in Step 3 is a constrained zonotope representation of the given H-Rep polytope $\mathcal{C}$. Since $\mathcal{C}$ is full-dimensional, $l<u$ and $\sigma<k$. Consequently,
$\left[\begin{array}[]{c}G_{C}\\ A_{C}\end{array}\right]=\left[\begin{array}[]{cc}\operatorname{diag}\left({\frac{u-l}{2}}\right)&0_{n\times M}\\ H\operatorname{diag}\left({\frac{u-l}{2}}\right)&\operatorname{diag}\left({\frac{\sigma-k}{2}}\right)\end{array}\right],$ (92)
is a lower triangular matrix with non-zero diagonal entries (and thereby, invertible). This completes the proof.
Proof of 5) Any Invertible representation satisfies
$\mathcal{C}$ =$=\left\{x\middle|\exists\xi\in\mathbb{R}^{{N_{C}}},\ {\|\xi\|}_{\infty}\leq 1,\ \left[G_{C};A_{C}\right]\xi=\left[x-c_{C};b_{C}\right]\right\}$
=$=\left\{x\middle|{\left\|{\left[G_{C};A_{C}\right]}^{-1}\left[x-c_{C};b_{C}\right]\right\|}_{\infty}\leq 1\right\}.$ (93)
We obtain (25) by rearranging terms in (93). ∎
A.2 Proof of Prop. 3
@@
$\mathcal{C}$ =$=\left\{x\middle|\sup_{\begin{subarray}{c}\nu\in\mathbb{R}^{n+{M_{C}}}\\ {\|[G_{C};A_{C}]^{\top}\nu\|}_{1}\leq 1\end{subarray}}\nu^{\top}[x-c_{C};b_{C}]\leq 1\right\}.$ (94)
For every $i\in\mathbb{N}_{[1:{N_{C}}]}$, define $\nu_{i}^{\ast}$ as the minimum-norm solution to $[G_{C};A_{C}]^{\top}\nu=e_{i}$. By Prop. 1.2, $\nu_{i}^{\ast}$ is available in closed-form, where
@@
for every $i\in\mathbb{N}_{[1:{N_{C}}]}$. Step 3 of Algo. 4 (see (48)) rescales $\nu_{i}^{\ast}$ to ensure that $[G_{C};A_{C}]^{\top}\nu_{i}^{\ast}$ has a unit $\ell_{1}$-norm to obtain $v_{i}^{\top}$. Thus, ±$\pm v_{i}^{\top}$ are feasible for (94) for every $i\in\mathbb{N}_{[1:{N_{C}}]}$, and using (94),
$\mathcal{C}$ $\subseteq\mathcal{P}\triangleq\left\{x\ \middle|\ {{\left\|V^{\top}[x-c_{C};b_{C}]\right\|}_{\infty}\leq 1}\right\},$ (95)
with $V=[\nu^{\ast}_{1};\ldots;\nu^{\ast}_{{N_{C}}}]$. Steps 45 of Algo. 4 define $\mathcal{P}=\{x\mid Hx\leq k\}$ with $H\in\mathbb{R}^{2{N_{C}}\times n}$ and $k\in\mathbb{R}^{2{N_{C}}}$. ∎
A.3 Proof of Prop. 5
From (93), an Invertible representation $\mathcal{C}$ may also be expressed as the following H-Rep polytope,
$\mathcal{C}$ =$=\left\{x\middle|\begin{array}[]{c}\forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\\ \delta e_{i}^{\top}{[G_{C};A_{C}]}^{-1}[x-c_{C};b_{C}]\leq 1\end{array}\right\}.$ (98)
Here, $v_{i}$ defined in Step 3 of Algo. 4 also simplifies to $e_{i}^{\top}{[G_{C};A_{C}]}^{-1}$ when $[G_{C};A_{C}]$ is invertible [21, Sec. 11.5]. Thus, Algo. 4 returns a H-Rep polytope $\mathcal{P}$ that coincides with (98) (thereby, equal to $\mathcal{C}$). The proof is completed with the observation that (98) is identical to (93). ∎
A.4 Proof of Prop. 6
Exactness of $\mathcal{M}^{+}$) As seen from Prop. 5, Algo. 4 computes an exact H-Rep polytope, given an Invertible representation $\mathcal{C}$. Since the rest of steps of Algo. 5 are exact, $\mathcal{M}^{+}=\mathcal{C}\ominus\mathcal{S}$.
Exactness of $\mathcal{M}^{-}$) For an Invertible representation $\mathcal{C}$, $\Gamma$ prescribed by Prop. 2 is
@@
by [21, Sec. 11.5]. We show that $\mathcal{M}=\mathcal{C}\ominus\mathcal{S}$ is an affine transformation of $\mathcal{C}_{L}$ defined in the proof of Thm. 1 (see (28)).
@@
=$\mathcal{M}=\left\{x\middle|\begin{array}[]{c}\ \forall s\in\mathcal{S}_{0},\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\\ \delta e_{i}^{\top}{[G_{C};A_{C}]}^{-1}[x+s+c_{S}-c_{C};b_{C}]\leq 1\end{array}\right\}$ (102)
@@
=$=\left\{x\ \middle|\ \begin{array}[]{c}\forall s\in\mathcal{S}_{0},\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\\ \delta e_{i}^{\top}\left({\Gamma x+\Gamma s+{[G_{C};A_{C}]}^{-1}[-c_{M^{-}};b_{C}]}\right)\leq 1\\ \end{array}\right\}$ (105)
@@
=$=\left\{x\ \middle|\ \begin{array}[]{c}\forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\\ \delta e_{i}^{\top}\left({\Gamma x+{[G_{C};A_{C}]}^{-1}[-c_{M^{-}};b_{C}]}\right)\leq 1-\rho_{\mathcal{S}_{0}}(\Gamma^{\top}e_{i})\\ \end{array}\right\}$ (108)
@@
=$=\left\{x\ \middle|\ \begin{array}[]{c}\xi\triangleq\Gamma x+{[G_{C};A_{C}]}^{-1}[-c_{M^{-}};b_{C}],\\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\xi\leq D_{ii}\\ \end{array}\right\}$ (111)
@@
=$=\left\{x\ \middle|\ \begin{array}[]{c}G_{C}\xi+c_{M^{-}}=x,\ A_{C}\xi=b_{C},\\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\xi\leq D_{ii}\\ \end{array}\right\}$ (114)
=$=\left\{G_{C}\xi+c_{M^{-}}\ \middle|\ A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ |e_{i}^{\top}\xi|\leq D_{ii}\right\}$
$=G_{C}\mathcal{C}_{L}+c_{M^{-}}=\mathcal{M}^{-}.$ (115)
@@
Here, (102) follows from (6d) and (98) in Sec. A.3, (105) follows from (99), (108) follows from encoding the condition for all $s\in\mathcal{S}_{0}$ using the support function of $\mathcal{S}_{0}$, (111) follows from the definition of $D_{ii}$ in (27), (114) follows from expressing $\xi$ as a solution of linear equations $[G_{C};A_{C}]\xi=[x-c_{M^{-}};b_{C}]$ with invertible $[G_{C};A_{C}]$, and (115) follows from the definition of $\mathcal{C}_{L}$ in (40) (see the proof of Thm. 1). Consequently, for an Invertible representation $\mathcal{C}$, Prop. 2 provides an exact characterization of $\mathcal{C}\ominus\mathcal{S}$. ∎
A.5 Proof of Prop. 7
@@
From [27, Thm. 2.1(iii)], $(\mathcal{U}\ominus\mathcal{V})\oplus\mathcal{V}\subseteq\mathcal{U}$ for any sets $\mathcal{U},\mathcal{V}$. Let a failure occur at some time $s\in\mathbb{N}$. Using (67) and [27, Thm. 2.1(iii)], for some $i\in\mathbb{N}_{[1:{L_{D}}]}$, $\hat{z}_{s|s}=x_{s}\in\mathcal{K}(s,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})\ominus\mathcal{E}_{\text{off-nom}}$, which implies that $\hat{z}_{s|s}\in\mathcal{K}(s,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})$. From Defn. 2, for every time step $k\in\mathbb{N}_{[s:s+T_{\text{safe}}]}$, there exists $\pi_{k}\in\Pi$ such that $u_{k|s}=\pi_{k}(\hat{z}_{k|s})\in\mathcal{U}_{\text{off-nom}}$ steers the state estimate according to the dynamics (66) to satisfy $\hat{z}_{k|s}\in\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$, despite the disturbance $\phi_{k|s}\in\Phi_{k}$. By (67) and [27, Thm. 2.1(iii)], $\hat{z}_{k|s}\in\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$ implies that $z_{k|s}\in\mathcal{H}_{i}^{\complement}$ (and thereby, $z_{k|s}\not\in\mathcal{D}$) for all $k\in\mathbb{N}_{[s:s+T_{\text{safe}}]}$, which completes the proof. ∎